# Certifying rings of integers in number fields

### v1

@@ 3.4. LPFW Certificate
There is a way of reducing irreducibility testing in $\mathbf{Z}[X]$ to primality testing in $\mathbf{Z}$: if a ’large enough’ integer $m$ is found such that $|f(m)|$ is prime, then $f$ can be shown to be irreducible. This test was proposed by Brillhart in (Brillhart, 1980). Abbott (Abbott, 2020), using a refined calculation that incorporates a known lower bound on the degrees of $f$’s factors, introduced the large prime factor witness (LPFW) certificate, which is similar to Brillhart’s, but works for a larger number of cases and often produces smaller prime witnesses. It relies on the following proposition, which we have formalized.
@@
249/130 := by rfl
With this, we formalized the LPFW certification scheme as:
@@
As before, irreducibility proofs for the factors of $f$ modulo primes can be handled as in Section 3.2. All remaining proofs included as fields in the above structure can be solved automatically using decide, rfl, and/or norm_num. We also formalized a variant with the trivial degree lower bound $d=1$, resulting in a shorter certificate.
We wrote a SageMath script that outputs a Lean proof of the irreducibility of $f$ by selecting the appropriate certificate from Section 3.3 or Section 3.4. As noted in Section 3.2, large polynomial degrees may cause issues with recursion depth or excesive heartbeats when checking the proofs in Lean.
In the search for a smaller prime witness $P$, it can be helpful to transform $f$. Möbius transformations  (Abbott, 2020), and more generally, Tschirnhaus transformation  (Cohen, 1993, Algorithm 6.3.4), can be used. Incorporating this into our proof generator is something we intend to pursue in the future.
@@ 4. Verifying rings of integers
As stated in the introduction, our main goal is to formally verify that a ring $\mathcal{O}$, with an explicitly given integral basis, equals the ring of integers of a number field $K$. Computation in $\mathcal{O}$ is key for both our proof and future applications. This raises two initial questions: how should we define $\mathcal{O}$ in Lean, and how can we compute in it?
@@ 4.1. Construction of the subalgebra
Defining $\mathcal{O}$ simply as the $R$-span of a given set gives a term of type Submodule R K. To make this into a subalgebra we also need proofs that it is closed under multiplication and contains $1$. In order to automate this construction starting with an $R$-basis, we created the SubalgebraBuilder.
Suppose that $R$ is a domain, $Q$ its fraction field, and $K$ the extension of $Q$ obtained by adjoining a root $\theta$ of a monic polynomial $T$ in $R[X]$ of degree $n$ (e.g. $R=\mathbf{Z}$, $Q=\mathbf{Q}$ and $K$ a number field). Suppose that $\{w_{1},\ldots,w_{n}\}$ is an $R$-basis for a subalgebra $\mathcal{O}$. Since $K\cong Q[X]/\langle T\rangle$, we can find a common denominator $d\in R\setminus\{0\}$ and polynomials $b_{i}\in R[X]$ of degree less than $n$ such that $w_{i}=\frac{1}{d}b_{i}(\theta)$.
As a module, $\mathcal{O}$ is the $R$-span of $\{w_{1},\ldots,w_{n}\}$. Showing that $\mathcal{O}$ is closed under multiplication amounts to proving that $w_{i}*w_{j}$ is in $\mathcal{O}$ for all $i$ and $j$. A way to certify this is to give the coefficients $a_{ijk}\in R$ such that $w_{i}*w_{j}=\sum_{k}a_{ijk}\cdot w_{k}$. Verifying this equality in $K$ is equivalent to proving that for all $i$ and $j$ there exists a polynomial $s_{ij}\in R[X]$ such that
@@
Thus, the elements $a_{ijk}$ and polynomials $s_{ij}$ act as a certificate for the closeness of $\mathcal{O}$ under multiplication. In fact, they provide a certificate for the precise representation of $w_{i}*w_{j}$ with respect to the basis $\{w_{1},\ldots,w_{n}\}$. Since multiplication is commutative, verifying this certificate requires checking $(n^{2}+n)/2$ identities in $R[X]$. Proving that $1$ is in $\mathcal{O}$ reduces to a single polynomial identity check.
@@
It takes as explicit parameters $R,Q,K$, the degree $n$, the monic polynomial $T$, and the list $L$ of coefficients of $T$. The fields of this structure include the matrix $B^{T}$ –with $B$ upper triangular containing the basis coefficients $c_{ij}$ as described above–, the common denominator $d$, the coefficients $a_{ijk}$, the certifying polynomials $s_{ij}$ in list form, as well as the proofs that $B$ is upper triangular with non-zero elements in the diagonal, that the identities in (1) hold (using list-based arithmetic) , and that $K$ is obtained by adjoining a root of $T$.
All of these proofs, except the latter, can be solved automatically using the decide and/or norm_num tactics. The latter statement is conveyed with the predicate IsAdjoinRoot K (map (algebraMap R Q) T)). Its proof will depend on the specific nature of $K$. If $K$ is defined as AdjoinRoot (map (algebraMap R Q) T), then this proof is obtained for free.
@@
is the corresponding $R$-subalgebra $\mathcal{O}$ of $K$ which, as a set, is the $R$-span of the set $\{w_{1},\ldots,w_{n}\}$, specified by $B$. We obtain a basis $(w_{i})_{i=1}^{n}$ for this subalgebra –indexed with Fin n– in basisOfBuilderLists T L A, with $w_{i}=\frac{1}{d}b_{i}(\theta)$.
@@ 4.2. Arithmetic in $\mathcal{O}$
More generally, let $S$ be an $R$-algebra which is finite and free as an $R$-module and has basis =$\mathcal{B}=\{w_{1},\ldots,w_{n}\}$. An element $x$ in $S$ can be uniquely written as $x=\sum_{i}c_{i}w_{i}$, and thus represented as a vector $(c_{1},\ldots,c_{n})$. We have that $S\cong R^{n}$ as $R$-modules. Furthermore, we can define a multiplication on $R^{n}$ that corresponds to the multiplication in $S$:
@@
We formally proved that, when $T$ is a times table for $S$, these functions on lists correspond to the arithmetic operations in $S$. Thus, identities in $S$ –involving additions and multiplications– can be proven by first translating them into their equivalent operations on lists. This method is used to prove identities in the subalgebra $\mathcal{O}$ of a number field for which we have a times table TimesTable (Fin n) $\mathcal{O}$. This verification on lists of integers can be done automatically by rfl or the decide tactic. A similar strategy is used for the $\mathbb{F}_{p}$-algebra $\mathcal{O}/p\mathcal{O}$ from Section 4.6.
@@ 4.3. Local maximality
Let $K$ be a number field of degree $n$ and $\mathcal{O}$ a subalgebra which is free and finite as a $\mathbf{Z}$-module. Then $\mathcal{O}$ is contained in $\mathcal{O}_{K}$, and we can compare them by considering the index $[\mathcal{O}_{K}:\mathcal{O}]$, typically defined as the size of the quotient group $\mathcal{O}_{K}/\mathcal{O}$. When $\mathcal{O}$ has rank $n$, then $[\mathcal{O}_{K}:\mathcal{O}]$ is a finite integer.
@@
[IsDomain R] [AddCommGroup M] [Module R M]. We formalized the above definition as:
@@
Subalgebra.toSubmodule ((Subalgebra.inclusion hm).range
@@
4.4. The Pohst-Zassenhaus theorem
For this section, let $R$ be a PID, $Q$ its fraction field, and $K$ an $R$-algebra. Consider two $R$-subalgebras of $K$, denoted by $\mathcal{O}$ and $\mathcal{O}_{m}$, which are free and finite as $R$-modules, have the same rank, and satisfy $\mathcal{O}\subseteq\mathcal{O}_{m}$. The primary case of interest to keep in mind is: $R=\mathbf{Z}$, $K$ is a number field, and $\mathcal{O}_{m}=\mathcal{O}_{K}$.
To construct a proof of $\mathcal{O}=\mathcal{O}_{m}$ we follow the strategy described in the previous section by establishing $\pi$-maximality for every prime $\pi$ of $R$. The main result we use is based on a theorem due to Pohst and Zassenhaus (Pohst and Zassenhaus, 1997, Lemma 5.53), which underpins the widely used Round-2 algorithm for computing rings of integers (Cohen, 1993, Algorithm 6.1.8).
@@ 4.4. The Pohst-Zassenhaus theorem
Let $\pi$ in $R$ be a prime and suppose that $\mathcal{O}$ and $\mathcal{O}_{m}$ have equal rank. If $\mathcal{O}=r(I_{\pi})$, then $\mathcal{O}$ is $\pi$-maximal.
This is a version of the Pohst-Zassenhaus theorem which is sufficient for our purposes. However, it is more commonly presented with $K$ being a finite dimensional and separable $Q$-algebra, and $\mathcal{O}_{m}$ the integral closure of $R$ in $K$. In that case, it is an if-and-only-if statement (Cohen, 1993), (Pohst and Zassenhaus, 1997).
@@
variable (hm : $\mathcal{O}$ $\mathcal{O}$m) $\pi$ : R} (hp : Prime $\pi$)
@@
( (Subalgebra.inclusion hm).range)
theorem order_piMaximal_of_order_eq_multiplierRing [Module.Free R $\mathcal{O}$m] [Module.Finite R $\mathcal{O}$m]
(heqr : Module.rank R $\mathcal{O}$ = Module.rank R $\mathcal{O}$m)
@@
As before, $\mathcal{O}*$ represents $\mathcal{O}$ as a submodule of $\mathcal{O}_{m}$. The term Ideal.span {algebraMap R O π} is the ideal $\pi\mathcal{O}$, with algebraMap R $\mathcal{O}$ π used to regard $\pi$ as an element of $\mathcal{O}$.
The proof strategy for this theorem, which we adapted from (Cohen, 1993), consists of defining an intermediate subalgebra $\mathcal{O}_{\pi}=\{x\in\mathcal{O}_{m}\mid\exists j\geq 0\text{ such that }\pi^{j}x\in\mathcal{O}\}$, which sits between $\mathcal{O}$ and $\mathcal{O}_{m}$ and is $\pi$-maximal by construction, and showing that $\mathcal{O}=\mathcal{O}_{\pi}$ whenever $\mathcal{O}=r(I_{\pi})$.
Using Theorem 4.1, the proof of $\pi$-maximality of $\mathcal{O}$ reduces to the verification of the equality $\mathcal{O}=r(I_{\pi})$. While an explicit description of $I_{\pi}$ may not be simple, in cases where $\mathcal{O}\cong R[X]/\langle T\rangle$ –with $T$ a monic polynomial– there is an explicit description of $I_{\pi}$ that leads to a simple criterion for $\pi$-maximality, called Dedekind criterion. The case where $\mathcal{O}$ is not necessarily of this form will be addressed in Section 4.6 for $R=\mathbf{Z}$.
@@ 4.5. Dedekind criterion
Let $g$ in $R[X]$ be a lift of the radical of $\bar{T}$. Set $h$ to be a lift of $\bar{T}/\bar{g}$ and define $f=(gh-T)/\pi$. We say that $T$ satisfies Dedekind criterion at $\pi$ if $\gcd(\bar{f},\bar{g},\bar{h})=1$.
If $K$ is a torsion-free $R$-algebra containing $\mathcal{O}$, then $T$ satisfying Dedekind criterion at $\pi$ implies $\mathcal{O}=r(I_{\pi})$. Combined with Theorem 4.1, this leads to the proposition:
@@
Let $K$ be torsion-free $R$-algebra containing $\mathcal{O}\cong R[X]/\langle T\rangle$. Let $\mathcal{O}_{m}$ be an $R$-subalgebra of $K$, containing $\mathcal{O}$ and with the same rank. If $T$ satisfies Dedekind criterion at $\pi$, then $\mathcal{O}$ is $\pi$-maximal (i.e. $\pi$ does not divide $[\mathcal{O}_{m}:\mathcal{O}]$).
To formalize the reduction modulo $\pi$ of polynomials as in Definition 4.2, perhaps the obvious choice would be to use the quotient map from R to R / (Ideal.span $\pi$}). However, for verifying Dedekind criterion we need to compute in $(R/\pi R)[X]$, and thus in $R/\pi R$. Hence, flexibility in the representation of $R/\pi R$ is important. For instance, we prefer to use polynomials over ZMod p rather than / (Ideal.span {p}) as we have decidable equality and efficient modular arithmetic in the former type but not in the latter.
For this reason, we model the quotient $R/\pi R$ as some type (F : Type*) with [Field F] instance, along with a ring homomorphism q : R →+* F which is surjective and its kernel equals Ideal.span $\pi$}. This ensures that $F\cong R/\pi R$. This approach of capturing the essential relationship between structures without fixing a specific construction is common in Mathlib as it allows for greater flexibility.
We formalized Dedekind criterion as follows:
@@
Formally, Proposition 4.3, adapted from (Cohen, 1993), reads:
@@
(hqker : RingHom.ker q = Ideal.span $\pi$}) (hmc : $\mathcal{O}$ $\mathcal{O}$m)
@@
[Module.Free R $\mathcal{O}$m] [Module.Finite R $\mathcal{O}$m]
@@
(heqr : Module.rank R $\mathcal{O}$= Module.rank R $\mathcal{O}$m)
@@
We also defined a specialized version of Dedekind criterion for the case $R=\mathbf{Z}$ and F = ZMod p. We created a structure, called CertificateDedekindCriterionLists, with parameters (T : ℤ[X]) (p : ℕ). Its fields include data certifying that $T$ satisfies Dedekind criterion at $p$, and proofs of the associated identities –involving additions and multiplications and using our list-based approach– which can be solved automatically by rfl or decide.
Dedekind criterion can still be useful even if $\mathcal{O}$ is not given by adjoining a root of a monic polynomial. If $\alpha\in\mathcal{O}$ has minimal polynomial $T\in R[X]$, then the $R$-subalgebra $R[\alpha]\subseteq\mathcal{O}$ is isomorphic to $R[X]/\langle T\rangle$. If $T$ satisfies Dedekind criterion at $\pi$ and $\mathcal{O}_{m}$ has rank equal to $\deg(T)$, then $R[\alpha]$, and hence $\mathcal{O}$, is $\pi$-maximal. Yet, for a number field $K$, the hope of using Dedekind criterion to prove $p$-maximality of its ring of integers at every prime $p$ soon fades. While this criterion always applies if $\mathcal{O}_{K}$ is monogenic, in some number fields the index $[\mathcal{O}_{K}:\mathbf{Z}[\alpha]]$ is divisible by a fixed prime for all $\alpha\in O_{K}$. An example was given by Dedekind (Dedekind, 1878).

### v2

@@ 4.4. LPFW Certificate
This section describes an alternative certificate for irreducibility of polynomials in $\mathbf{Z}[X]$ which is more widely applicable. We discuss its formalization in Lean, and a script that can automatically generate a Lean proof of irreducibility based on these certificates.
There is a way of reducing irreducibility testing in $\mathbf{Z}[X]$ to primality testing in $\mathbf{Z}$: if a ‘large enough’ integer $m$ is found such that $|f(m)|$ is prime, then $f$ can be shown to be irreducible. This test was proposed by Brillhart in (Brillhart, 1980). Abbott (Abbott, 2020), using a refined calculation that incorporates a known lower bound on the degrees of $f$’s factors, introduced the large prime factor witness (LPFW) certificate, which is similar to Brillhart’s, but works for a larger number of cases and often produces smaller prime witnesses. It relies on the following proposition, which we have formalized.
@@
249/130 := by decide!
Recall from Section 3 that decide! performs kernel reduction to prove decidable propositions.
We formalized the LPFW certification scheme as:
@@
As before, irreducibility proofs for the factors of $f$ modulo primes can be handled as in Section 4.2. All remaining proofs included as fields in the above structure can be solved automatically using decide, decide!, and/or norm_num. We also formalized a variant with the trivial degree lower bound $d=1$, resulting in a shorter certificate.
We wrote a SageMath script that outputs a Lean proof of the irreducibility of $f$ by selecting the appropriate certificate from Section 4.3 or Section 4.4. As noted in Section 4.2, large polynomial degrees may cause issues with recursion depth or excesive heartbeats when checking the proofs in Lean. This approach of generating Lean code from SageMath worked well for our applications. However, exploring a tighter integration, such as calling SageMath directly from Lean, could be an interesting direction for future work.
In the search for a smaller prime witness $P$, it can be helpful to transform $f$. Möbius transformations (Abbott, 2020), and more generally, Tschirnhaus transformation (Cohen, 1993, Algorithm 6.3.4), can be used. Incorporating this into our proof generator is something we also intend to pursue in the future.
@@ 5. Verifying Rings of Integers
As stated in the introduction, our main goal is to formally verify that a ring $\mathcal{O}$, with an explicitly given basis and contained inside a number field $K$, equals –in the set-theoretic sense– the ring of integers of $K$. Computation in $\mathcal{O}$ is key for both our proof and future applications. This raises two initial questions: how should we define $\mathcal{O}$ in Lean, and how can we compute in it?
@@ 5.1. Construction of the Subalgebra
The goal of this section is to describe the construction of a ring $\mathcal{O}\subseteq K$ in Lean, given by an explicit basis.
@@
Defining $\mathcal{O}$ simply as the $R$-span of a given set gives a term of type Submodule R K. To make this into a subalgebra we also need proofs that it is closed under multiplication and contains $1$. In order to automate this construction in Lean starting with an $R$-basis, we created the SubalgebraBuilder.
To describe how it works, let us further specialize our setup and fix some notation. Suppose that $R$ is a domain, $Q$ its fraction field, and $K$ the extension of $Q$ obtained by adjoining a root $\theta$ of a monic polynomial $T$ in $R[X]$ of degree $n$ (e.g. $R=\mathbf{Z}$, $Q=\mathbf{Q}$ and $K$ a number field). Suppose that $\{w_{1},\ldots,w_{n}\}$ is an $R$-basis for a subalgebra $\mathcal{O}$ of $K$. Since $K\cong Q[X]/\langle T\rangle$, we can find a common denominator $d\in R\setminus\{0\}$ and polynomials $b_{i}\in R[X]$ of degree less than $n$ such that $w_{i}=\frac{1}{d}b_{i}(\theta)$.
As a module, $\mathcal{O}$ is the $R$-span of $\{w_{1},\ldots,w_{n}\}$. However, as previously remarked, in order to give it a type Subalgebra R K instead of just Submodule R K, we need to convince Lean that $\mathcal{O}$ is closed under multiplication and contains $1$. To prove closure under multiplication it suffices to show that $w_{i}*w_{j}$ is in $\mathcal{O}$ for all $i$ and $j$. A way to certify this is to give the coefficients $a_{ijk}\in R$ such that $w_{i}*w_{j}=\sum_{k}a_{ijk}\cdot w_{k}$. Verifying this equality in $K$ is equivalent to proving that for all $i$ and $j$ there exists a polynomial $s_{ij}\in R[X]$ such that
@@
Thus, the elements $a_{ijk}$ and polynomials $s_{ij}$ act as a certificate for the closure of $\mathcal{O}$ under multiplication. In fact, they provide a certificate for the precise representation of $w_{i}*w_{j}$ with respect to the basis $\{w_{1},\ldots,w_{n}\}$. Since multiplication is commutative, verifying this certificate requires checking $(n^{2}+n)/2$ identities in $R[X]$. Proving that $1$ is in $\mathcal{O}$ reduces to a single polynomial identity check.
@@
It takes as explicit parameters $R,Q,K$, the degree $n$, the monic polynomial $T$, and the list $L$ of coefficients of $T$. We follow the practice in Mathlib to encode the structure of inclusion of $R$ into its fraction field $Q$, and $Q$ into the ring $K$ using the Algebra and IsFractionRing typeclasses. The transitive inclusion of $R$ into $K$ is witnessed by [Algebra R K] [IsScalarTower R Q K] instances; the latter asserts the triangle of inclusions $R\to Q\to K$ commutes. Browning and Lutz (Browning and Lutz, 2022) explain Mathlib’s approach to extensions (of fields) in depth, and Wieser (Wieser, 2021) covers scalar actions in Mathlib more generally.
The fields of the structure include the matrix $B^{T}$ –with $B$ upper triangular containing the basis coefficients $c_{ij}$ as described above–, the common denominator $d$, the coefficients $a_{ijk}$, the certifying polynomials $s_{ij}$ in list form, as well as the proofs that $B$ is upper triangular with non-zero elements in the diagonal, that the identities in (1) hold (using list-based arithmetic), and that $K$ is obtained by adjoining a root of $T$. All of these proofs, except the latter, can be solved automatically using the decide, decide!, and/or norm_num tactics. The latter statement is conveyed with the predicate IsAdjoinRoot K (map (algebraMap R Q) T). Its proof will depend on the specific nature of $K$. If $K$ is defined as AdjoinRoot (map (algebraMap R Q) T), then this proof is obtained for free.
@@
is the corresponding $R$-subalgebra $\mathcal{O}$ of $K$ which, as a set, is the $R$-span of $\{w_{1},\ldots,w_{n}\}$, specified by $B$. We obtain a basis $(w_{i})_{i=1}^{n}$ for this subalgebra –indexed with Fin n– in basisOfBuilderLists T L A, with $w_{i}=\frac{1}{d}b_{i}(\theta)$.
@@ 5.2. Arithmetic in $\mathcal{O}$
More generally, let $S$ be an $R$-algebra which is finite and free as an $R$-module and has basis =$\mathcal{B}=\{w_{1},\ldots,w_{n}\}$. An element $x$ in $S$ can be uniquely written as $x=\sum_{i}c_{i}w_{i}$, and thus represented as a vector $(c_{1},\ldots,c_{n})\in R^{n}$. We have that $S\cong R^{n}$ as $R$-modules. Furthermore, we can define a multiplication on $R^{n}$ that corresponds to the multiplication in $S$:
@@
We formally proved that, when $T$ is a times table for $S$, these functions on lists correspond to the arithmetic operations in $S$. Thus, identities in $S$ –involving additions and multiplications– can be proven by first translating them into their equivalent operations on lists. This method is used to prove identities in the subalgebra $\mathcal{O}$ of a number field for which we have a times table TimesTable (Fin n) $\mathcal{O}$. This verification on lists of integers can be done automatically by a tactic such as decide!. A similar strategy is used for the $\mathbb{F}_{p}$-algebra $\mathcal{O}/p\mathcal{O}$ from Section 5.6.
@@ 5.3. Local Maximality
Here, we present a local notion of maximality which serves as a tool to prove an equality between subalgebras.
Let $K$ be a number field of degree $n$ and $\mathcal{O}$ a subalgebra which is free and finite as a $\mathbf{Z}$-module. Then $\mathcal{O}$ is contained in $\mathcal{O}_{K}$, and we can compare them by considering the index $[\mathcal{O}_{K}:\mathcal{O}]$, typically defined as the cardinality of the quotient $\mathcal{O}_{K}/\mathcal{O}$. When $\mathcal{O}$ has rank $n$, then $[\mathcal{O}_{K}:\mathcal{O}]$ is a finite integer.
@@
[IsDomain R] [AddCommGroup M] [Module R M]. We formalized the above notion of index as:
@@
Subalgebra.toSubmodule (Subalgebra.inclusion hm).range
@@
5.4. The Pohst–Zassenhaus Theorem
In this section, we introduce a crucial theorem that can be used to establish local maximality and forms the basis for the certificate discussed in Section 5.6.
Let $R$ be a PID, $Q$ its fraction field, and $K$ an $R$-algebra. Consider two $R$-subalgebras of $K$, denoted by $\mathcal{O}$ and $\mathcal{O}^{\prime}$, which are free and finite as $R$-modules, have the same rank, and satisfy $\mathcal{O}\subseteq\mathcal{O}^{\prime}$. The primary case of interest to keep in mind is: $R=\mathbf{Z}$, $K$ is a number field, and $\mathcal{O}^{\prime}=\mathcal{O}_{K}$.
To construct a proof of $\mathcal{O}=\mathcal{O}^{\prime}$ we follow the strategy described in the previous section by establishing $\pi$-maximality for every prime $\pi$ of $R$. The main result we use is based on a theorem due to Pohst and Zassenhaus (Pohst and Zassenhaus, 1997, Lemma 5.53), which underpins the widely used Round-2 algorithm for computing rings of integers (Cohen, 1993, Algorithm 6.1.8).
@@ 5.4. The Pohst–Zassenhaus Theorem
Let $\pi$ in $R$ be a prime and suppose that $\mathcal{O}$ and $\mathcal{O}^{\prime}$ have equal rank. If $\mathcal{O}=r(I_{\pi})$, then $\mathcal{O}$ is $\pi$-maximal.
This is a version of the Pohst–Zassenhaus theorem which is sufficient for our purposes. However, it is more commonly presented with $K$ being a finite dimensional and separable $Q$-algebra, and $\mathcal{O}^{\prime}$ the integral closure of $R$ in $K$. In that case, it is an if-and-only-if statement (Cohen, 1993), (Pohst and Zassenhaus, 1997).
@@
variable (hm : $\mathcal{O}$ $\mathcal{O}$’) $\pi$ : R} (hp : Prime $\pi$)
@@
(Subalgebra.inclusion hm).range
theorem order_piMaximal_of_order_eq_multiplierRing [Module.Free R $\mathcal{O}$’] [Module.Finite R $\mathcal{O}$’]
(heqr : Module.rank R $\mathcal{O}$ = Module.rank R $\mathcal{O}$’)
@@
As before, $\mathcal{O}*$ represents $\mathcal{O}$ as a submodule of $\mathcal{O}^{\prime}$. The term Ideal.span {algebraMap R O π} is the ideal $\pi\mathcal{O}$, with algebraMap R $\mathcal{O}$ π used to regard $\pi$ as an element of $\mathcal{O}$.
The proof strategy for this theorem, which we adapted from (Cohen, 1993), consists of defining an intermediate subalgebra $\mathcal{O}_{\pi}=\{x\in\mathcal{O}^{\prime}\mid\exists j\geq 0\text{ such that }\pi^{j}x\in\mathcal{O}\}$, which sits between $\mathcal{O}$ and $\mathcal{O}^{\prime}$ and is $\pi$-maximal by construction, and showing that $\mathcal{O}=\mathcal{O}_{\pi}$ whenever $\mathcal{O}=r(I_{\pi})$.
Using Theorem 5.1, the proof of $\pi$-maximality of $\mathcal{O}$ reduces to the verification of the equality $\mathcal{O}=r(I_{\pi})$. While an explicit description of $I_{\pi}$ may not be simple, in cases where $\mathcal{O}\cong R[X]/\langle T\rangle$ –with $T$ a monic polynomial– there is an explicit description of $I_{\pi}$ that leads to a simple criterion for $\pi$-maximality, called the Dedekind criterion. The case where $\mathcal{O}$ is not necessarily of this form will be addressed in Section 5.6 for $R=\mathbf{Z}$.
@@ 5.5. Dedekind Criterion
We present a practical criterion involving only polynomial arithmetic that provides a simple way of establishing $\pi$-maximality in certain cases.
@@
Let $g$ in $R[X]$ be a lift of the radical of $\bar{T}$. Set $h$ to be a lift of $\bar{T}/\bar{g}$ and define $f=(gh-T)/\pi$. We say that $T$ satisfies the Dedekind criterion at $\pi$ if $\gcd(\bar{f},\bar{g},\bar{h})=1$.
If $K$ is a torsion-free $R$-algebra containing $\mathcal{O}$, then $T$ satisfying the Dedekind criterion at $\pi$ implies $\mathcal{O}=r(I_{\pi})$. Combined with Theorem 5.1, this leads to the proposition:
@@
Let $K$ be torsion-free $R$-algebra containing $\mathcal{O}\cong R[X]/\langle T\rangle$. Let $\mathcal{O}^{\prime}$ be an $R$-subalgebra of $K$, containing $\mathcal{O}$ and with the same rank. If $T$ satisfies the Dedekind criterion at $\pi$, then $\mathcal{O}$ is $\pi$-maximal (i.e. $\pi$ does not divide $[\mathcal{O}^{\prime}:\mathcal{O}]$).
To formalize the reduction modulo $\pi$ of polynomials as in Definition 5.2, perhaps the obvious choice would be to use the quotient map from R to R / (Ideal.span $\pi$}). However, for verifying the Dedekind criterion we need to compute in $(R/\pi R)[X]$, and thus in $R/\pi R$. Hence, flexibility in the representation of $R/\pi R$ is important. For instance, we prefer to use polynomials over ZMod p rather than / (Ideal.span {p}) as we have decidable equality and efficient modular arithmetic in the former type but not in the latter.
For this reason, we model the quotient $R/\pi R$ as some type (F : Type*) with [Field F] instance, along with a ring homomorphism q : R →+* F which is surjective and its kernel equals Ideal.span $\pi$}. This ensures that $F\cong R/\pi R$. This approach of capturing the essential relationship between structures without fixing a specific construction is common in Mathlib as it allows for greater flexibility. In this context, a proof transfer framework such as (Cohen et al., 2024) could also be useful for automating the translation of results involving different representations of $R/\pi R$, though we did not explore this further.
We formalized the Dedekind criterion as follows:
@@
Observe that the discussed assumptions on q : R →+* F are not needed to write down the previous definition. However, they do appear in the following theorem, which is the formal version of Proposition 5.3, adapted from (Cohen, 1993):
@@
(hqker : RingHom.ker q = Ideal.span $\pi$}) (hmc : $\mathcal{O}$ $\mathcal{O}$’)
@@
[Module.Free R $\mathcal{O}$’] [Module.Finite R $\mathcal{O}$’]
@@
(heqr : Module.rank R $\mathcal{O}$= Module.rank R $\mathcal{O}$’)
@@
We also defined a specialized version of the Dedekind criterion for the case $R=\mathbf{Z}$ and F = ZMod p. We created a structure, called CertificateDedekindCriterionLists, with parameters (T : ℤ[X]) (p : ℕ). Its fields include data certifying that $T$ satisfies the Dedekind criterion at $p$, and proofs of the associated identities –involving additions and multiplications and using our list-based approach– which can be solved automatically by rfl or decide.
The Dedekind criterion can still be useful even if $\mathcal{O}$ is not given by adjoining a root of a monic polynomial. If $\alpha\in\mathcal{O}$ has minimal polynomial $T\in R[X]$, then the $R$-subalgebra $R[\alpha]\subseteq\mathcal{O}$ is isomorphic to $R[X]/\langle T\rangle$. If $T$ satisfies the Dedekind criterion at $\pi$ and $\mathcal{O}^{\prime}$ has rank equal to $\deg(T)$, then $R[\alpha]$, and hence $\mathcal{O}$, is $\pi$-maximal. Yet, for a number field $K$, the hope of using the Dedekind criterion to prove $p$-maximality of its ring of integers at every prime $p$ soon fades. While this criterion always applies if $\mathcal{O}_{K}$ is monogenic, in some number fields the index $[\mathcal{O}_{K}:\mathbf{Z}[\alpha]]$ is divisible by a fixed prime for all $\alpha\in\mathcal{O}_{K}$. Dedekind gave an example of this (Dedekind, 1878).